16. The Bank of New England is concerned about the amount of debt being accrued by customers using its credit cards. The Board of directors decided to install an expensive monitoring system if the mean for all of the banks customers is greater than $2000. The bank randomly selected 100 credit card holders and determined the amounts they charged. For this sample group the mean is $2177 and standard deviation is $843. Use 0.025 level of significance to test the claim that the mean amount charged is greater than $2000. Based on the results of testing, will the monitoring system be implemented? (7points). agr Or ric
Answer #1

Null Hypothesis H0: Mean amount charged = $2000

Alternative Hypothesis Ha: Mean amount charged > $2000

As, the alternative hypothesis has the greater than sign, it is a right tail test.

Level of significance = 0.025

Standard error of sample mean = 843/V100 = 84.3

Test statistic , t = (Observed mean - Hypothesized mean) / Standard error = (2177 - 2000) / 84.3 = 2.10

Degree of freedom, df = 100 - 1= 99

For df = 99, P[t > 2.1] = 0.0191

As, p-value is less than the 0.025 level of significance, we reject null hypothesis and conclude there is statistical significance evidence that the Mean amount charged > $2000.

As, we have concluded that mean amount charged is greater than $2000, the monitoring system should be implemented.